determine the slope and y intercept of the graph of 12+3y=-2x
Answer by Mathtut(3670)

write in slope intercept form which is y=mx+b.where m is the slope and b is the y intercept
:
3y=-2x+12
:
y=(-2/3)x+4........m=-2/3 and y intercept is 4:(0,4)